OpenStudy (anonymous):

@iGreen solve using substitution please 2x+5y=2 -3x-y=-3

OpenStudy (welshfella):

from second equation find y in terms of x then replace the y in the first equation with this expression in x sole for x

OpenStudy (anonymous):

nevermind sorry I meant this equation

OpenStudy (welshfella):

* solve

OpenStudy (anonymous):

2x-3y=4 x=-4y-9

OpenStudy (welshfella):

ok the easiest way is to replace the x in first equation by -4y-9:- 2(-4y-9) - 3y = 4 now solve this for y

OpenStudy (welshfella):

can you do that?

OpenStudy (anonymous):

-8y-18 what do I do with the -3y=4?

OpenStudy (welshfella):

-8y - 18 - 3y = 4 add like terms -11y - 18 = 4 add 18 to both sides -11y = 22 divide both sides by -11 y = -2

OpenStudy (welshfella):

now to find the value of x you substitute y = -2 into the first equation and solve for x

OpenStudy (anonymous):

I got -1? is that right?

OpenStudy (welshfella):

2x - 3(-2) = 4 2x = + 6 = 4 2x = -2 x = -1 yes

OpenStudy (welshfella):

* 2x + 6 = 4

OpenStudy (welshfella):

good work

OpenStudy (anonymous):

thanks please stay I need more help.

OpenStudy (welshfella):

ok

OpenStudy (welshfella):

but i have to go in 10 minutes

OpenStudy (anonymous):

sorry computer froze x+y=15 4x+3y=38

OpenStudy (welshfella):

do you want to solve this by substitution also?

OpenStudy (anonymous):

yes

OpenStudy (welshfella):

x + y = 15 subtract x from both sides y = 15 - x now plug this into the second equation:- 4x + 3(15 - x) = 38 4x + 45 - 3x = 38 (4x - 3x = x so:- x + 45 = 38 now subtract 45 form both sides to get value of x

OpenStudy (welshfella):

then plug this value of x into the equation x = y = 15 to get value of y

OpenStudy (welshfella):

* x + y = 15

OpenStudy (anonymous):

x=7?

OpenStudy (welshfella):

38 - 45 = -7

OpenStudy (welshfella):

x + 45 = 38 x + 45 - 45 = 38 - 45 x = -7

OpenStudy (anonymous):

okay got that so plug -7 into x+y=15?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

-7 + y = 15

OpenStudy (welshfella):

gotta go in a few minutes

OpenStudy (anonymous):

okay thanks so much is the answer 22 though?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

yw